OpenStudy (anonymous):

-3/4+1/6u=3/2 what is u?? Please help!!!!

OpenStudy (anonymous):

u =27/2

OpenStudy (anonymous):

so 1/6u = 9/4 4=54u u=27/2

OpenStudy (anonymous):

thanks!!